# fev-bench: A Realistic Benchmark for Time Series Forecasting

### v1

@@ Abstract
Benchmark quality is critical for meaningful evaluation and sustained progress in time series forecasting, particularly given the recent rise of pretrained models. Existing benchmarks often have narrow domain coverage or overlook important real-world settings, such as tasks with covariates. Additionally, their aggregation procedures often lack statistical rigor, making it unclear whether observed performance differences reflect true improvements or random variation. Many benchmarks also fail to provide infrastructure for consistent evaluation or are too rigid to integrate into existing pipelines. To address these gaps, we propose fev-bench, a benchmark comprising 100 forecasting tasks across seven domains, including 46 tasks with covariates. Supporting the benchmark, we introduce fev, a lightweight Python library for benchmarking forecasting models that emphasizes reproducibility and seamless integration with existing workflows. Using fev, fev-bench employs principled aggregation methods with bootstrapped confidence intervals to report model performance along two complementary dimensions: win rates and skill scores. We report results on fev-bench for various pretrained, statistical and baseline models, and identify promising directions for future research.
@@ 1 Introduction
Pretrained time series forecasting models are transforming forecasting practice. They often deliver more accurate forecasts than traditional methods (Aksu et al., 2024), while enabling zero-shot inference that simplifies production use and lowers barrier to entry for practitioners (Cohen et al., 2025).
Advances in pretrained forecasting models are largely assessed empirically on benchmarks, making benchmark quality essential for continued progress. Oversights in forecasting benchmarks propagate directly to the model development. For example, most general forecasting benchmarks completely ignore covariates despite their prevalence in real-world applications (Bojer & Meldgaard, 2021; Arango et al., 2025). As a result, the majority of pretrained models lack covariate support, limiting their effectiveness in domains like retail where promotional data and pricing information are essential for accurate demand forecasting (Fildes et al., 2022).
Beyond task coverage, existing benchmarks often lack statistical rigor. Most studies report single-number summaries, leaving it unclear whether improvements reflect true advances or random variation. Small gains may vanish or even reverse with minor benchmark changes (Roque et al., 2025). This undermines the reliability of conclusions and can mislead researchers and practitioners about which models perform better.
Finally, benchmark infrastructure presents additional barriers to progress and reproducibility. Many benchmarks provide only standalone datasets without evaluation code, leading to inconsistent implementations across studies that make results incomparable (Hewamalage et al., 2023). When infrastructure exists, it often consists of monolithic systems bundling models, datasets, and evaluation logic with extensive dependencies that become unmaintainable over time. These rigid systems prevent researchers from extending benchmarks to new domains or integrating evaluation into existing workflows, limiting their practical utility and lifespan.
To address these challenges, we make the following three contributions:
  • New benchmark. We introduce fev-bench, a forecast evaluation benchmark containing 100 tasks spanning 7 real-world application domains. Our benchmark addresses a key gap in existing work by including 46 tasks with covariates alongside both univariate and multivariate forecasting scenarios, better reflecting real-world forecasting use cases.
@@
  • Evaluation package. We introduce fev 111github.com/autogluon/fev, a lightweight Python package for forecasting evaluation that introduces minimal dependencies while remaining compatible with popular forecasting libraries. The package focuses on reproducibility and extensibility, enabling researchers to easily build and share new benchmarks and the corresponding results.
@@ 2 Preliminaries
Problem definition. The multivariate time series forecasting problem can be formally stated as follows. We are given a collection $\{{\bm{y}}_{n,1:T}\}_{n=1}^{N}$ of $N$ multivariate time series. For $n=1,\dots,N$ and $t=1,\dots,T$, let ${\bm{y}}_{n,t}=(y_{n,d,t})_{d=1}^{D}\in\mathbb{R}^{D}$ denote the $D$-dimensional observation vector for series $n$ at time $t$, where the special case $D=1$ corresponds to univariate forecasting. The goal of multivariate forecasting is to predict the future $H$ values ${\bm{y}}_{n,T+1:T+H}$ for each series $n$, where $H$ is the forecast horizon. Each time series may be accompanied by covariates $\mathbf{X}_{n,1:T+H}$, which include (i) static covariates that do not vary over time (e.g., item or location identifiers), (ii) past-only dynamic covariates observed up to time $T$ (e.g., past measurements of related variables), and (iii) known dynamic covariates available for all time steps $1,\dots,T+H$ (e.g., holiday indicators or planned interventions).
In the most general form, the aim is to model the conditional distribution
$p\big({\bm{y}}_{n,T+1:T+H}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H}\big).$ (1)
While this full distributional modeling provides the richest information, it is common in practice to produce point forecasts like the conditional mean or median of each $y_{n,d,t}$. Alternatively, it is often sufficient to estimate predictive quantiles of the conditional distributions $p(y_{n,d,t}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H})$. We denote by ${\mathcal{Q}}\subset(0,1)$ the set of quantile levels of interest and aim to produce forecasts $\hat{y}^{(q)}_{n,d,t}$ such that $\Pr(y_{n,d,t}\leq\hat{y}^{(q)}_{n,d,t})=q$ for all $q\in{\mathcal{Q}}$.
@@
Tasks. A task specifies a concrete forecasting problem. It consists of a dataset together with all parameters that define how forecasts are produced and evaluated, including the forecast horizon $H$, the evaluation cutoff dates, which columns serve as targets or covariates, and the evaluation metric. A single dataset can yield multiple distinct tasks by varying these parameters. The choice of evaluation metric is integral to task definition because different metrics can correspond to different optimal forecasts; combining conflicting metrics within a single task creates ambiguity about the intended goal (Kolassa, 2020).
Rolling evaluation. Each task is evaluated using a rolling-origin evaluation protocol with $W$ windows (Hyndman & Athanasopoulos, 2018). Let $\tau_{1}<\tau_{2}<\dots<\tau_{W}$ denote the evaluation cutoffs. At each window $w\in\{1,\dots,W\}$, the model receives all observations up to $\tau_{w}$ as input and is asked to produce $H$-step forecasts. Advancing the cutoff creates a sequence of forecast–target pairs, so that a single task produces $W$ evaluation windows. This setup mimics real-world deployment and yields a more robust estimate of model performance over time, especially for datasets consisting of few series.
Aggregation. While tasks define individual evaluation problems, benchmarks also require a method to aggregate results across tasks, enabling us to answer questions like “Is model A more accurate than model B overall?”. The aggregation method directly affects the reliability and interpretability of benchmark results.
We discuss task construction in Section 3, aggregation procedures in Section 4, and the software package supporting the benchmark in Section 5.
@@ 3 Task definitions
We introduce fev-bench (Forecast EValuation Benchmark), a comprehensive benchmark designed to address the limitations of existing forecasting evaluation frameworks. Unlike previous benchmarks that focus on narrow domains or do not provide the evaluation infrastructure, fev-bench provides broad coverage across real-world applications while ensuring reproducible and statistically sound comparisons.
fev-bench comprises 100 forecasting tasks with complete specifications provided in Appendix A. This section explains our design choices that guided its construction.
@@ 3.1 Datasets and tasks
Our goal is to construct a general benchmark that is representative of various real-world forecasting applications. To achieve this, the benchmark must cover different domains, frequencies, horizons, and time series characteristics. We consider both univariate and multivariate forecasting problems, with covariates covering dynamic (both past-only and known) and static variables. We evaluate both point and probabilistic forecasting performance.
Datasets. We begin by sourcing time series datasets from established collections including Monash repository (Godahewa et al., 2021), GIFT-Eval (Aksu et al., 2024), and BOOM (Cohen et al., 2025). However, these collections lack datasets with covariates. To address this critical gap, we extend our benchmark by incorporating public datasets from Kaggle (Bojer & Meldgaard, 2021) and domain-specific repositories (Wang et al., 2023; Lago et al., 2021; Arango et al., 2025).
This curation process yields 96 unique time series datasets.222huggingface.co/datasets/autogluon/fev_datasets We construct 100 forecasting tasks from these datasets through different target column and covariate selections. Among these tasks, 30 include known dynamic covariates, 24 include past dynamic covariates, and 19 include static covariates. These categories are non-exclusive (a task may include multiple types of covariates) and both univariate and multivariate tasks can have associated covariates. Tables 2 and 3 compare our benchmark’s coverage against GIFT-Eval, the most comprehensive existing general forecasting benchmark. fev-bench provides substantially more datasets across various domains and frequencies.
Forecast horizons. Many existing benchmarks reuse identical datasets with different horizons (Zeng et al., 2023; Aksu et al., 2024), which creates correlated tasks that provide limited additional insights into model performance. While fev-bench includes both short and long-horizon tasks, we deliberately avoid horizon duplication within datasets. Instead, we select horizons that reflect domain-appropriate forecasting needs, such as 168 steps for hourly energy demand or 30 steps for daily retail sales.
Rolling evaluation. We use rolling window evaluation to balance computational cost with statistical reliability. The number of windows $W$ is determined by dataset size: up to 20 windows for datasets with fewer than 10 series, up to 10 windows for datasets with 10–2000 series, and 1 window for larger datasets. The actual value of $W$ is constrained by available data length, ensuring at least $(2\times H+1)$ past observations before the first evaluation window. We aggregate the results across rolling windows with arithmetic mean.
@@ 3.2 Evaluation metrics
Each of the 100 tasks in our benchmark evaluates both point and probabilistic forecast accuracy using complementary metrics.
We evaluate point forecast accuracy using Mean Absolute Scaled Error (MASE), following most existing benchmarks (Aksu et al., 2024; Godahewa et al., 2021).
@@
where $\hat{y}_{n,d,t}$ is the point forecast and $a_{n,d}$ is the historical seasonal error of series $n$ along dimension $d$, defined as
$a_{n,d}=\frac{1}{T-m}\sum_{t=m+1}^{T}|y_{n,d,t}-y_{n,d,t-m}|.$ (3)
Here, $m$ is the seasonal period determined by the data frequency (e.g., m=12m{=}12 for monthly data with yearly seasonality). MASE offers several advantages: it is scale-free, balances contributions across series with different magnitudes, handles trends well, and remains robust when the forecast horizon contains zeros (Hyndman & Koehler, 2006).
We evaluate the probabilistic forecast accuracy using the Scaled Quantile Loss (SQL) computed on quantile levels ${\mathcal{Q}}=\{0.1,0.2,...,0.9\}$:
@@
where $\hat{y}^{(q)}_{n,d,t}$ is the quantile forecast at level $q$, $a_{n,d}$ is the historical seasonal error (Equation 3), and $\rho_{q}(\cdot,\cdot)$ is the quantile loss
@@ 3.3 Representative subset of the tasks
In addition to the full fev-bench benchmark consisting of 100 tasks, we provide fev-bench-mini, a curated subset of 20 tasks. These tasks are chosen to capture the diversity of covariates, dimensionalities, domains, and horizons present in the full benchmark, while being small enough to enable rapid iteration and reduced computational cost. fev-bench-mini approximates the relative performance ranking observed on the full benchmark, making it suitable for model development, ablation studies, and resource-constrained evaluations. More details on fev-bench-mini are provided in Appendix E.
@@ 4.1 Marginal performance
The primary goal of any benchmark is to rank models by their average performance. We employ two complementary aggregation methods that capture different aspects of model quality.
@@ 4.3 Significance of performance differences
We report bootstrap confidence intervals only for the pairwise statistics ($W_{jk},S_{jk}$), as these directly answer the question of interest: “Does model $j$ consistently outperform model $k$ under different benchmark compositions?”. Confidence intervals for the marginal statistics ($W_{j},S_{j}$) instead describe the variability of each model’s average score in isolation, ignoring the correlations between models.
@@ 5 Infrastructure
Comprehensive task coverage and principled evaluation are important, but standardized infrastructure is equally vital for the relevance and longevity of a benchmark. This includes code to define tasks, run evaluations, and aggregate results in a consistent manner.
@@ 5.1 Motivation
Existing forecasting benchmarks usually fall into one of two categories: standalone datasets without supporting infrastructure (Godahewa et al., 2021; Zeng et al., 2023) and end-to-end systems which bundle models, datasets and forecasting tasks (Qiu et al., 2024; Aksu et al., 2024).
Benchmarks with standalone datasets provide no guarantees that the results obtained by different users are comparable. Two users may have evaluated on distinct forecasting tasks even when referring to the same dataset. This could be due to differences in the forecast horizon, the forecast start date, or the evaluation metric implementation. Even with identical task specifications, differences in the aggregation strategy can dramatically changing the final conclusion.
While standalone datasets are problematic due to ambiguity, end-to-end systems with models, datasets and tasks are often impractical due to their rigidity. These systems usually come with lots of dependencies and assumptions, which makes extending or integrating them into existing workflows difficult. The model implementations become stale over time as the maintenance overhead is often too high. Restrictions on commonly-used libraries such as torch, numpy, and pandas lead to “dependency hell”. Differences in library versions also make the evaluation non-transparent: For example, a different pandas version may change how time series frequencies are inferred, which affects the seasonal period $m$, which in turn alters the computation of MASE.
@@
To address these limitations, we introduce fev, a lightweight library that provides essential benchmarking functionality without unnecessary constraints or bloated dependencies. Its core features include task definition, data loading and splitting, prediction scoring, and result aggregation.
fev only depends on Hugging Face datasets (Lhoest et al., 2021) and pydantic (Colvin et al., 2025) libraries for input validation and does not fix versions of commonly-used packages like torch or numpy, allowing hassle-free integration in existing model pipelines. Using datasets enables effortless data loading, avoiding custom file formats or complex data processing pipelines. Moreover, it makes the library future-proof and it can support use cases such as multimodal forecasting (with text and image features) out of the box.
@@
fev does not include model implementations, which have the potential of becoming outdated. Instead, it provides adapters which transform data into formats expected by popular forecasting packages like GluonTS (Alexandrov et al., 2020), darts (Herzen et al., 2022), Nixtla libraries (Garza et al., 2022; Olivares et al., 2022), AutoGluon (Shchur et al., 2023), and sktime (Löning et al., 2019).
5.2 fev API
The fev library is built around three main constructs:
  • EvaluationWindow — a single train–test split of time series data, defined by a cutoff index or date. This is the smallest unit on which metrics can be computed.
  • Task — a complete specification of a forecasting problem. It includes the dataset, forecast horizon, initial cutoff, covariates (past-only, future-known, and static), target columns used for evaluation, evaluation metric(s), and any metric-specific parameters such as the seasonal period used for MASE. A Task may include one or more EvaluationWindows.
  • Benchmark — a collection of Tasks.
Benchmarks in fev can be defined using YAML files, allowing users to easily construct custom benchmarks. Each task produces an evaluation summary containing not only the metric values but also the full task specification, ensuring that the results are unambiguous. These summaries make it straightforward to compare results and immediately identify any differences in setup. Additionally, fev provides utilities for aggregating results across tasks in a benchmark, as described in Section 4.

### v2

@@ Abstract
Benchmark quality is critical for meaningful evaluation and sustained progress in time series forecasting, particularly with the rise of pretrained models. Existing benchmarks often have limited domain coverage or overlook real-world settings such as tasks with covariates. Their aggregation procedures frequently lack statistical rigor, making it unclear whether observed performance differences reflect true improvements or random variation. Many benchmarks lack consistent evaluation infrastructure or are too rigid for integration into existing pipelines. To address these gaps, we propose fev-bench, a benchmark of 100 forecasting tasks across seven domains, including 46 with covariates. Supporting the benchmark, we introduce fev, a lightweight Python library for forecasting evaluation emphasizing reproducibility and integration with existing workflows. Using fev, fev-bench employs principled aggregation with bootstrapped confidence intervals to report performance along two dimensions: win rates and skill scores. We report results on fev-bench for pretrained, statistical, and baseline models and identify promising future research directions.
@@ 1 Introduction
Pretrained time series forecasting models are transforming forecasting practice. They often outperform traditional methods (Aksu et al., 2024) while enabling zero-shot inference that simplifies production use and lowers the barrier to entry (Cohen et al., 2025).
Advances in pretrained forecasting models are primarily evaluated on benchmarks, making benchmark quality critical for progress. Shortcomings in existing benchmarks directly shape model development; for example, most general benchmarks ignore covariates despite their prevalence in real-world applications (Bojer & Meldgaard, 2021; Arango et al., 2025). Consequently, most pretrained models lack covariate support, limiting their effectiveness in domains such as retail, where promotional and pricing data are essential for accurate demand forecasting (Fildes et al., 2022).
Beyond task coverage, existing benchmarks often lack statistical rigor. Most studies report single-number summaries, making it unclear whether improvements reflect true advances or random variation. Small gains may vanish or even reverse under minor benchmark changes (Roque et al., 2025), undermining the reliability of conclusions about which models perform better.
Finally, benchmark infrastructure poses additional barriers to progress and reproducibility. Many benchmarks provide only standalone datasets without evaluation code, leading to inconsistent implementations and incomparable results (Hewamalage et al., 2023). When infrastructure exists, it often consists of monolithic systems that bundle models, datasets, and evaluation logic with extensive dependencies, becoming unmaintainable over time. This rigidity prevents extension to new domains or integration into existing workflows, limiting practical utility and longevity.
To address these challenges, we make three contributions:
  • New benchmark. We introduce fev-bench, a forecast evaluation benchmark containing 100 tasks spanning 7 real-world application domains.111huggingface.co/spaces/autogluon/fev-bench Our benchmark addresses a key gap in existing work by including 46 tasks with covariates alongside both univariate and multivariate forecasting scenarios, better reflecting real-world forecasting use cases.
@@
  • Evaluation package. We introduce fev, 222github.com/autogluon/fev a lightweight Python package for forecasting evaluation that introduces minimal dependencies while remaining compatible with popular forecasting libraries. The package focuses on reproducibility and extensibility, enabling researchers to easily build and share new benchmarks and the corresponding results.
@@ 2 Preliminaries
Problem definition. The multivariate time series forecasting problem can be formally stated as follows. We are given a collection $\{{\bm{y}}_{n,1:T}\}_{n=1}^{N}$ of $N$ multivariate time series. For $n=1,\dots,N$ and $t=1,\dots,T$, let ${\bm{y}}_{n,t}=(y_{n,d,t})_{d=1}^{D}\in\mathbb{R}^{D}$ denote the $D$-dimensional observation vector for series $n$ at time $t$, where $D=1$ corresponds to univariate forecasting. The goal is to predict the future $H$ values ${\bm{y}}_{n,T+1:T+H}$ for each series $n$, where $H$ is the forecast horizon. Each time series may be accompanied by covariates $\mathbf{X}_{n,1:T+H}$, including (i) static covariates that do not vary over time (e.g., item or location identifiers), (ii) past-only dynamic covariates observed up to time $T$ (e.g., past related series), and (iii) known dynamic covariates available for all time steps $1,\dots,T+H$ (e.g., holiday indicators).
In the most general form, the goal is to model the conditional distribution $\smash{p\big({\bm{y}}_{n,T+1:T+H}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H}\big)}$. While full distributional modeling provides the richest information, it is common in practice to produce point forecasts such as conditional means or medians. Alternatively, many applications estimate predictive quantiles $\smash{p(y_{n,d,t}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H})}$. We denote by ${\mathcal{Q}}\subset(0,1)$ the set of quantile levels and produce forecasts $\smash{\hat{y}^{(q)}_{n,d,t}}$ satisfying $\smash{\Pr(y_{n,d,t}\leq\hat{y}^{(q)}_{n,d,t})=q}$ for all $q\in{\mathcal{Q}}$.
@@
Tasks. A task specifies a concrete forecasting problem. It consists of a dataset and the parameters that define how forecasts are produced and evaluated, including the forecast horizon $H$, evaluation cutoffs, target and covariate columns, and the evaluation metric. A single dataset can yield multiple tasks by varying these parameters. The choice of evaluation metric is integral to task definition, as different metrics correspond to different optimal forecasts; combining conflicting metrics within a single task creates ambiguity about the intended goal (Kolassa, 2020).
Rolling evaluation. Each task is evaluated using a rolling-origin evaluation protocol with $W$ windows (Hyndman & Athanasopoulos, 2018). Let $\tau_{1}<\tau_{2}<\dots<\tau_{W}$ denote the evaluation cutoffs. At each window $w\in\{1,\dots,W\}$, the model receives observations up to $\tau_{w}$ and produces $H$-step forecasts. Advancing the cutoff yields a sequence of forecast–target pairs, resulting in $W$ evaluation windows per task. This setup mimics real-world deployment and provides a more robust estimate of performance over time, especially for datasets with few series.
Aggregation. While tasks define individual evaluation problems, benchmarks must aggregate results across tasks to answer questions such as “Is model A more accurate than model B overall?”. The choice of aggregation directly affects the reliability and interpretability of benchmark results.
We discuss task definitions in Sec. 3, aggregation methods in Sec. 4, and the accompanying software package in Sec. 5.
@@ 3 Task definitions
We introduce fev-bench (Forecast EValuation Benchmark), a benchmark designed to address limitations of existing forecasting evaluation frameworks. Unlike prior benchmarks that focus on narrow domains or lack evaluation infrastructure, fev-bench offers broad coverage of real-world applications while enabling reproducible and statistically sound comparisons. fev-bench comprises 100 forecasting tasks, with complete specifications in Appendix A. This section describes the design choices guiding its construction.
@@ 3.1 Datasets and tasks
Our goal is to construct a general benchmark representative of real-world forecasting applications across domains, frequencies, horizons, and time series characteristics. We consider both univariate and multivariate forecasting problems, with covariates covering dynamic (both past-only and known) and static variables. We evaluate both point and probabilistic forecasting performance.
Datasets. We source time series datasets from established collections including the Monash repository (Godahewa et al., 2021), GIFT-Eval (Aksu et al., 2024), and BOOM (Cohen et al., 2025). However, these collections lack datasets with covariates. To address this gap, we extend the benchmark with public datasets from Kaggle (Bojer & Meldgaard, 2021) and domain-specific repositories (Wang et al., 2023; Lago et al., 2021; Arango et al., 2025).
This curation process yields 96 unique time series datasets.333huggingface.co/datasets/autogluon/fev_datasets We construct 100 forecasting tasks from these datasets using different target column and covariate selections. Among these tasks, 30 include known dynamic covariates, 24 include past dynamic covariates, and 19 include static covariates. These categories are non-exclusive, and both univariate and multivariate tasks may include covariates. Tables 2 and 3 compare our benchmark’s coverage against GIFT-Eval, the most comprehensive existing general forecasting benchmark. fev-bench provides substantially broader coverage across domains and frequencies.
Forecast horizons. Many existing benchmarks reuse the same datasets with different horizons (Zeng et al., 2023; Aksu et al., 2024), creating correlated tasks with limited additional insight. While fev-bench includes both short- and long-horizon tasks, we avoid horizon duplication within datasets. Instead, we select horizons that reflect domain-specific forecasting needs, such as 168 steps for hourly energy demand or 30 steps for daily retail sales.
Rolling evaluation. We use rolling window evaluation to balance computational cost and statistical reliability. The number of windows $W$ depends on dataset size: up to 20 windows for datasets with fewer than 10 series, up to 10 windows for datasets with 10–2000 series, and one window for larger datasets. The value of $W$ is further constrained by data availability, requiring at least $(2\times H+1)$ past observations before the first evaluation window. Results are aggregated across windows using the arithmetic mean.
@@ 3.2 Evaluation metrics
Each of the 100 benchmark tasks evaluates point and probabilistic forecast accuracy using complementary metrics. We evaluate point forecast accuracy using Mean Absolute Scaled Error (MASE), following existing benchmarks (Aksu et al., 2024; Godahewa et al., 2021).
@@
where the error for each series $n$ and dimension $d$ is normalized by the historical seasonal error $\smash{a_{n,d}=\frac{1}{T-m}\sum_{t=m+1}^{T}|y_{n,d,t}-y_{n,d,t-m}|}$. Here $\hat{y}_{n,d,t}$ is the point forecast and $m$ is the seasonal period based on the data frequency (e.g., m=12m{=}12 for monthly data). MASE offers several advantages: it is scale-free, balances contributions across series with different magnitudes, handles trends well, and remains robust when the forecast horizon contains zeros (Hyndman & Koehler, 2006).
We use Scaled Quantile Loss (SQL) computed on levels ${\mathcal{Q}}=\{0.1,0.2,...,0.9\}$ for probabilistic forecast accuracy:
@@
where $\hat{y}^{(q)}_{n,d,t}$ is the quantile forecast at level $q$, $a_{n,d}$ is the historical seasonal error as in MASE, and $\rho_{q}(\cdot,\cdot)$ is the quantile loss
@@ 3.3 Representative subset of the tasks
In addition to the full fev-bench benchmark consisting of 100 tasks, we provide fev-bench-mini, a curated subset of 20 tasks (Appendix E). These tasks are chosen to capture the diversity of covariates, dimensionalities, domains, and horizons present in the full benchmark, while being small enough to enable rapid iteration and reduced computational cost. fev-bench-mini approximates the relative performance ranking observed on the full benchmark, making it suitable for model development and ablation studies.
@@ 4.1 Marginal performance
The primary goal of any benchmark is to rank models by average performance. We use two complementary aggregation methods that capture different aspects of model quality.
@@ 4.3 Significance of performance differences
We report bootstrap confidence intervals only for the pairwise statistics ($W_{jk},S_{jk}$), as these directly answer the question of interest: “Does model $j$ consistently outperform model $k$ under different benchmark compositions?” Equivalently, the intervals define paired null-hypothesis tests over tasks, with null hypotheses $W_{jk}=0.5$ (no systematic advantage) and $S_{jk}=0$ (no average error reduction). Confidence intervals for the marginal statistics ($W_{j},S_{j}$) instead describe the variability of each model’s average score in isolation and ignore correlations between models.
@@ 5 Infrastructure
Comprehensive task coverage and principled evaluation are essential, but standardized infrastructure is equally important for benchmark relevance and longevity. This includes code for task definition, evaluation, and result aggregation.
@@ 5.1 Motivation
Existing forecasting benchmarks typically fall into one of two categories: standalone datasets without supporting infrastructure (Godahewa et al., 2021; Zeng et al., 2023), and end-to-end systems that bundle models, datasets, and forecasting tasks (Qiu et al., 2024; Aksu et al., 2024).
Benchmarks consisting only of datasets provide no guarantees that results obtained by different users are comparable. Users may evaluate different forecasting tasks even when referring to the same dataset, due to differences in horizons, cutoffs, metrics, or aggregation strategies.
End-to-end systems address this ambiguity but are often impractical due to their rigidity. They impose many dependencies and assumptions, making extension or integration into existing workflows difficult. Model implementations also tend to become outdated as maintenance costs grow. Version constraints on commonly used libraries such as torch, numpy, and pandas can further reduce transparency. For example, changes in pandas may affect frequency inference, altering the seasonal period $m$ and MASE computation.
5.2 fev Python library
@@
To address these limitations, we introduce fev, a lightweight library that provides essential benchmarking functionality without unnecessary constraints. Its core features include task definition, data loading and splitting, prediction scoring, and result aggregation. fev depends only on Hugging Face datasets (Lhoest et al., 2021) and pydantic (Colvin et al., 2025) for input validation, and does not fix versions of commonly used packages such as torch or numpy. This enables seamless integration into existing pipelines. Using datasets also simplifies data loading and supports extensions such as multimodal forecasting.
@@
fev deliberately excludes model implementations, which are prone to becoming outdated. Instead, it provides adapters that convert data into formats compatible with popular forecasting libraries including GluonTS (Alexandrov et al., 2020), darts (Herzen et al., 2022), StatsForecast (Garza et al., 2022), AutoGluon (Shchur et al., 2023), and sktime (Löning et al., 2019).
Library API. The fev library is built around three main constructs. An EvaluationWindow is a single train–test split defined by a cutoff. A Task specifies a forecasting problem, including the dataset, horizon, covariates, targets, and evaluation metric, and may include multiple evaluation windows. Finally, a Benchmark is a collection of tasks. Benchmarks in fev can be defined using YAML files. Each task produces an evaluation summary that includes both metric values and the full task specification, ensuring unambiguous and comparable results. fev also provides utilities for aggregating results across tasks, as described in Sec. 4.